# Scalable Fine-tuning from Multiple Data Sources: A First-Order Approximation Approach

### v1

@@ 3.3 Accelerating subset selection for LMs
GradEx-DS: In data selection, one would like to pick a data subset from a collection of raw data. To apply our technique, one can preprocess the raw data by clustering it into $n$ groups. Then, use the above procedures to choose a subset from the $n$ groups.
@@
One can write down the intermediate calculations plus carry bits. We thus generate a synthetic addition set by generating input pairs of length $5$ between 00 and $9$. We create $10$ groups; five are correct, while the rest involve randomly generated output digits. Then, we aim to separate the noisy groups from the correctly-labeled groups.
@@
Figure 2: Illustrating the separation between correctly-labeled and noisy examples, using estimated values of $T$ in equation (1), compared to several existing measures. Our measure from $T$ (rightmost) leads to a much more clear separation of correct vs. noisy examples compared to three measures.
@@ 3.4 Comparison with prior approaches
In particular, our algorithm reduces $m$ model fine-tuning runs to a single meta-model training. As mentioned earlier, the additional estimation stage incurs very little overhead; after dimension reduction, solving each logistic regression problem takes less than 2 seconds per task subset. This estimation stage takes less than 10% of the total computation cost in our overall procedure. As for the meta-training stage, our algorithm uses comparable computational costs to existing data selection methods such as Xie et al. (2023) and Xia et al. (2024). Table 2 summarizes this comparison, including the number of forward passes required by each approach.
@@
Table 2: Summary of runtime complexity between our algorithm and existing solutions for subset selection, as a function of the number of data sources $n$. We describe the constants in terms of forward passes each method takes. For every method, the number of backward passes equals the forward passes. Here, $\alpha$ denotes the size of subsets sampled in random ensembles. Note that the number of forward passes is for one training step; for the calculation, see Appendix C.
4 Experiments
We now validate GradEx and its use cases across various datasets and models, focusing on the following key questions. Does the estimation procedure accurately approximate the true model fine-tuning losses? How much computational cost does it save relative to classical subset selection methods that require repeated model fine-tuning? How effective are the subset selection methods using the estimation results in downstream evaluation?
Our experiments show that GradEx approximates fine-tuned model losses within 1% error, tested on five different LMs including Llama-3-8B. GradEx reduces the number of FLOPs by up to 43$\times$ and GPU hours by 46$\times$ compared to conventional subset selection. Next, we evaluate GradEx for instruction tuning and chain-of-thought fine-tuning. In both cases, our algorithm performs on par with conventional subset selection that uses full fine-tuing results, while incurring less than 0.5% computation costs. Our algorithm outperforms existing selection methods Xie et al. (2023); Xia et al. (2024) on ToxiGen by 3.8% and TruthfulQA by 2.4%, while using comparable computation costs.
4.1 Experimental setup
Our algorithm is broadly applicable to estimating model fine-tuning performances; We focus on instruction tuning and chain-of-thought fine-tuning in this section. For instruction tuning, given a set of source tasks and a target task, our goal is to select source tasks relevant to the target task. For chain-of-thought fine-tuning, we fine-tune an LM to generate chain-of-thought reasoning steps and answers to answer a question. Several explanations are possible, while some of them are incorrect. We aim to select explanations pertinent to the reasoning task using subset selection.
Table 3: We report the relative RSS between $\hat{f}(S)$ and $f(S)$, measured on Alpaca and StrategyQA. For measuring speedup, we report the ratio of the number of FLOPs required for computing $f(S)$ divided by GradEx. The speedup remains the same across different models for the same dataset, since the speed-up stems from a reduced number of forward/backward passes.
@@ 4.1 Experimental setup
For chain-of-thought fine-tuning, we use CommonSenseQA Talmor et al. (2019) and StrategyQA Geva et al. (2021). The chain-of-thought explanations are generated with a GPT-3 175B model. We sample 10% of the data for evaluating $f(S)$.
We set LoRA as the base protocol. We adjust the rank parameter of LoRA between 16, 32, 64, and 128. For chain-of-thought fine-tuning, we partition the data into $100$ groups using spectral clustering on the gradient similarity matrix. For random ensembles, we sample $1000$ subsets, each containing 75% of all tasks, to ensure the $T$ scores (cf. equation (1)) have converged.
4.2 Approximating model fine-tuning losses
We now assess the accuracy of GradEx. We measure the relative error between the estimated $\hat{f}(S)$ and the true $f(S)$:
@@ 4.2 Approximating model fine-tuning losses
We obtain $f(S)$ by fine-tuning a pretrained LM on the samples of $S$ (along with the target task). We measure computation cost through the total number of FLoating-point OPerations (FLOPs), and the number of GPU hours (measured on a desktop with three Nvidia RTX6000 GPUs).
@@
Figure 3: Number of GPU hours as the number of model parameters between our estimation approach (left, 3(a)) and full fine-tuning (right, 3(b)). We estimate the full fine-tuning cost by fine-tuning on randomly sampled 100 subsets of tasks.
(a) Task selection for instruction tuning ($n=1,729$)
(b) Data selection for chain-of-thought fine-tuning ($n=100$)
Figure 4: Illustration of the tradeoff between the number of FLOPs (computation cost) and test error rate, measured on our methods and six baseline methods. Our approach delivers comparable downstream performance to conventional subset selection methods, while using less than 0.5% of total computation. Furthermore, our approach achieves a 3.1% performance improvement over existing data selection methods using comparable computation costs. The comparison of GPU hours is qualitatively similar and can be found in Appendix D.3.
On Alpaca, we find that GradEx-FS uses 117 GPU hours for Llama-3-8B; this is 17.6$\times$ less computation compared to running forward selection with full fine-tuning. As for random ensembles, GradEx-RE uses 120 GPU hours for Llama-3-8B and 43.3$\times$ less computation compared to full fine-tuning. We also observe qualitatively similar results when measured on StrategyQA: GradEx achieves 30.5$\times$ and 44.8$\times$ less computation compared to full fine-tuning in forward selection and random ensembles, respectively.
We note that the speedup remains consistent across different models when applied to the same dataset. Figure 3 illustrates the number of GPU hours used by GradEx when running random ensembles on five LMs on Alpaca (left, 4(a)) vs. full fine-tuning (right, 4(b)).
@@
Reducing overfitting in meta-training: Recall that in the meta-training stage, we apply multitask training to the combined samples of all the tasks. We conduct a preliminary experiment, where we use a sharpness-reduced training algorithm (i.e., sharpness-aware minimization Foret et al. (2021); Zhang et al. (2024)) for multitask training (in place of SGD). We find that this can reduce the approximation error (relative to full fine-tuning results) by 23%. Further applying multitask and meta-learning techniques to improve GradEx would be an interesting question for future work.
@@ 4.3 Downstream evaluation for fine-tuning
Baselines: We compare our algorithms to forward selection (FS) and random ensemble (RE) with full fine-tuning. Additionally, we compare with four baseline methods, including fine-tuning a single model on all tasks with LoRA (MTL), Data Selection with Importance Resampling (DSIR) Xie et al. (2023), Data Efficient Fine-Tuning with cross-task nearest neighbors clustering based on feature similarity (DEFT) Ivison et al. (2023), and Low-rank gradiEnt Similarity Search (LESS) Xia et al. (2024). For each baseline, the performance is reported based on fine-tuning a pretrained LM with the same amount of trainable parameters on the selected subset of tasks.
Evaluation metrics: For instruction tuning, we follow the protocol of Wang et al. (2023a) to evaluate the toxicity and truthfulness of fine-tuned models in ToxiGen Hartvigsen et al. (2022) and TruthfulQA Lin et al. (2022), respectively. For ToxiGen, we measure the percentage of toxic outputs generated by the model. For TruthfulQA, we evaluate model accuracy for identifying truthful statements. For chain-of-thought fine-tuning, we measure the accuracy of two reasoning tasks.
@@ 4.3.1 Results for instruction tuning
First, we compare GradEx-FS with using the predictions from the pretrained LM directly, using the features from the LM in a regression problem, and fine-tuning on all tasks (MTL). We find that GradEx-FS outperforms each of these three methods by 9%, 11%, and 8% on average, respectively.
@@ 4.3.2 Results for chain-of-thought fine-tuning
First, GradEx-FS outperforms directly using the pretrained model for making predictions, feature transfer from the pretrained model, and fine-tuning on all tasks by 7%, 16%, and 5% (on average), respectively. Second, even with the estimation we see similar downstream performance to forward selection and random ensemble (with full fine-tuning), while using only 3% and 0.5% of the total computation costs. Lastly, both GradEx-FS and GradEx-RE can outperform the four baseline selection methods by 2.4% and 3.5% (on average), respectively.
As for data selection, we first cluster the samples into $n$ groups. We vary $n$ between $50$, $100$, $200$, and $400$ and we also evaluate our approach without the clustering step on StrategyQA. We find that using $n=100$ groups yields the best performance, and indeed outperforms not using clustering by 1.4%. Therefore, we set $n=100$ in all the data selection experiments.
@@ 5 Related Work
Data modeling: There is a growing line of work on understanding the role of individual samples in deep networks and large models. Wettig et al. (2024) select pre-training data by training a rater model to evaluate four data quality criteria that align with human intuitions. Thrush et al. (2024) select data based on the losses and perplexities of existing LLMs on pretraining texts. In contrast, this work focuses on task-relatedness in fine-tuning language models. This is a simpler problem compared to pre-training, as the number of tasks is relatively smaller. A potential adaptation could be splitting pre-training into multiple phases and applying our techniques to develop a curriculum for pre-training. Li et al. (2023b) introduce a surrogate modeling approach to identify subsets of source tasks that benefit a target task. They show that a linear surrogate model is particularly helpful for identifying negative transfers, including higher-order negative transfers from a subset of source tasks to the target task. This higher-order transfer corresponds to a boosting procedure that is helpful Li et al. (2023a). Li et al. (2024a) introduce a tree-structured surrogate model for finding compositions of data augmentations. Our approach is also related to the work of Park et al. (2023) on using random sampling for data attribution. The difference is that we utilize the linearization property of LLMs for fine-tuning close to the initialization.
@@ 6 Conclusion
Acknowlegement: Thanks to the anonymous referees and the action editor for their constructive feedback. The work of D. Li, Z. Zhang, and H. Zhang is in part supported by NSF award IIS-2412008.
@@ Appendix C Ommited Materials for Section 3
2% $6_{\pm 0.6}\times 10^{-4}$ $6_{\pm 0.5}\times 10^{-5}$ $8_{\pm 1.1}\times 10^{-4}$ $3_{\pm 0.4}\times 10^{-5}$
4% $1_{\pm 0.2}\times 10^{-3}$ $3_{\pm 0.3}\times 10^{-4}$ $2_{\pm 0.4}\times 10^{-3}$ $1_{\pm 0.3}\times 10^{-4}$
6% $2_{\pm 0.1}\times 10^{-3}$ $6_{\pm 0.6}\times 10^{-4}$ $3_{\pm 0.4}\times 10^{-4}$ $3_{\pm 0.7}\times 10^{-4}$
8% $3_{\pm 0.3}\times 10^{-3}$ $1_{\pm 0.2}\times 10^{-3}$ $3_{\pm 0.7}\times 10^{-4}$ $4_{\pm 0.9}\times 10^{-4}$
10% $4_{\pm 0.2}\times 10^{-2}$ $5_{\pm 0.8}\times 10^{-3}$ $7_{\pm 0.6}\times 10^{-2}$ $5_{\pm 0.8}\times 10^{-3}$
@@ D.3 Comparison of GPU hours
# FLOPs ($\downarrow$)
@@ D.5 Complete pseudocode for subset selection procedures
Input: Training datasets of $n$ data sources; a multitask fine-tuning algorithm $f$
@@
2:for Step $i=1,2,\ldots,n$ do
3:     for Each dataset not selected in current subset $j\in[n]/S_{\text{cur}}$  do
4:         Add dataset $j$ to current subset: $S_{j}\leftarrow S_{\text{cur}}\bigcup\{j\}$
5:         Evaluate $\hat{f}(S_{j})$ using GradEx
@@
7:     Choose the dataset with the highest value function $j^{*}\leftarrow\arg\max_{j\in[n]/S_{\text{cur}}}\hat{f}(S<{j})$
8:     Add the dataset to the current subset: $S_{\text{cur}}\leftarrow S_{\text{cur}}+j^{*}$, if $\hat{f}(S_{j})>\hat{f}(S_{\text{cur}})$. Otherwise, stop searching
@@
3:     Evaluate $\hat{f}(S_{k})$ using GradEx
@@
5:Estimate the random ensemble score for each dataset $i$ as the value function averaged over subsets that include $i$:
@@
6:Select a subset of datasets by thresholding the scores: $S^{\star}=\{i\ |\ \hat{T}_{i}>\gamma,\ \forall\ i=1,2,\ldots,n\}$

### v2

@@ 3.3 Accelerating subset selection for LMs
GradEx-DS: In data selection, one would like to pick a data subset from a collection of raw data. To apply our technique, one can preprocess the raw data by clustering it into $n$ groups. Then, use the above procedures to choose a subset from the $n$ groups of data samples.
@@
One can write down the intermediate calculations plus carry bits. We thus generate a synthetic addition set by generating input pairs of length $5$ between 00 and $9$. We create $20$ groups; ten are correct, while the rest involve randomly generated output digits. Then, we aim to separate the noisy groups from the correctly-labeled groups.
@@
Figure 2: Illustrating the separation between correctly-labeled and noisy examples, using estimated values of $T$ in equation (1), compared to several existing measures. Our measure from $T$ (rightmost) leads to a much more clear separation of correct vs. noisy examples compared to the three measures. In particular, the second and third figures use gradient-based feature similarity, and our loss-based selection procedure leads to a more accurate separation than these measures.
@@ 3.4 Comparison with prior approaches
In particular, our algorithm reduces $m$ model fine-tuning runs to a single meta-model training. As mentioned earlier, the additional estimation stage incurs very little overhead; after dimension reduction, solving each logistic regression problem takes less than 2 seconds per task subset. This estimation stage takes less than 10% of the total computation cost in our overall procedure. As for the meta-training stage, our algorithm uses comparable computational costs to existing data selection methods such as Xie et al. (2023) and Xia et al. (2024). Table 2 summarizes this comparison, including the number of forward passes required by each approach. Note that the number of forward passes is for one training step. For every method, the number of backward passes equals the forward passes. For the calculation of each method, see Appendix C for the details.
Table 2: Summary of runtime complexity between our algorithm and existing solutions for subset selection, as a function of the number of data sources $n$. We describe the constants in terms of forward passes each method takes. Here, $\alpha$ denotes the size of subsets sampled in random ensembles.
@@
Table 3: We report the relative RSS between $\hat{f}(S)$ and $f(S)$, measured on Alpaca and StrategyQA. For measuring speedup, we report the ratio of the number of FLOPs required for computing $f(S)$ divided by GradEx. The speedup remains the same across different models for the same dataset since the speed-up stems from a reduced number of forward and backward passes.
@@
4 Experiments
We now validate GradEx and its use cases across various datasets and models, focusing on the following key questions. Does the estimation procedure accurately approximate the true model fine-tuning losses? How much computational cost does it save relative to classical subset selection methods that require repeated model fine-tuning? How effective are the selection methods based on estimated results in downstream evaluation?
Our experiments show that GradEx approximates fine-tuned model losses within 1% error, tested on five different LMs, including Llama-3-8B. GradEx reduces the number of FLOPs by up to 43$\times$ and GPU hours by 46$\times$ compared to conventional subset selection. Next, we evaluate GradEx for instruction tuning and chain-of-thought fine-tuning. In both cases, our algorithm performs on par with conventional subset selection that uses true fine-tuning results while incurring less than 0.5% computation costs. Our algorithm outperforms existing selection methods Xie et al. (2023); Xia et al. (2024) on ToxiGen by 3.8% and TruthfulQA by 2.4%, while using comparable computation costs.
4.1 Experimental setup
Our algorithm is broadly applicable to estimating LM fine-tuning performances. In this section, we focus on instruction tuning and chain-of-thought fine-tuning. For the former, given a set of source tasks and a target task, we aim to select source tasks relevant to the target task. For the latter, we fine-tune an LM to generate chain-of-thought reasoning steps and the answer to a question. Several explanations are possible, while some of them are incorrect. We aim to select explanations pertinent to the reasoning task using subset selection.
@@ 4.1 Experimental setup
For chain-of-thought fine-tuning, we use CommonSenseQA Talmor et al. (2019) and StrategyQA Geva et al. (2021) datasets. The chain-of-thought explanations are generated with a GPT-3 175B model. We sample 10% of the data for evaluating $f(S)$. We partition the data into $100$ groups by applying spectral clustering to a matrix of gradient cosine similarity scores between every pair of samples. Each group refers to one task. In the end, we have $100$ tasks from the $100$ groups in both the CommonSenseQA and StrategyQA datasets.
We set LoRA as the base protocol. We adjust the rank parameter of LoRA between 16, 32, 64, and 128. For random ensembles, we sample $1000$ subsets, each containing 75% of all tasks, to ensure the $T$ scores (cf. equation (1)) have converged.
4.2 Approximating language model fine-tuning losses
We now assess the accuracy of GradEx. We measure the relative error between the estimated $\hat{f}(S)$ and the true value of $f(S)$ (which can be obtained by running full fine-tuning) as the averaged error over $m$ subsets:
@@ 4.2 Approximating language model fine-tuning losses
More specifically, we obtain $f(S)$ by fine-tuning a pretrained LM on the samples of $S$ along with the samples from the target task. We measure computation cost through the total number of floating-point operations (FLOPs) and the number of GPU hours (measured on a desktop with three Nvidia RTX6000 GPUs).
@@
(a) Task selection for instruction tuning ($n=1,729$)
(b) Data selection for chain-of-thought fine-tuning ($n=100$)
Figure 3: Illustration of the tradeoff between the number of FLOPs (computation cost) and test error rate, measured on our methods and six baseline methods. Our approach delivers comparable downstream performance to conventional subset selection methods, while using less than 0.5% of total computation. Furthermore, our approach achieves a 3.8% performance improvement over existing data selection methods using comparable computation costs. In particular, MTL refers to the standard approach of training one model on the data of all tasks, and we can see that our approach can indeed reduce the error rates compared to MTL. The comparison of GPU hours is qualitatively similar and can be found in Appendix D.3.
On Alpaca, we find that GradEx-FS uses 117 GPU hours for Llama-3-8B; This is 17.6$\times$ less computation compared to running forward selection with true fine-tuning. As for random ensembles, GradEx-RE uses 120 GPU hours for Llama-3-8B and 43.3$\times$ less computation compared to true fine-tuning. We also observe qualitatively similar results when measured on StrategyQA: GradEx achieves 30.5$\times$ and 44.8$\times$ less computation compared to true fine-tuning in the forward selection and random ensembles, respectively.
@@
Figure 4: Number of GPU hours as the number of model parameters between our estimation approach (left, Figure 4(a)) and full fine-tuning (right, Figure 4(b)). We estimate the full fine-tuning cost by fine-tuning on randomly sampled 100 subsets of tasks.
We note that the speedup remains consistent across different models when applied to the same dataset. Figure 4 illustrates the number of GPU hours used by GradEx when running our approach on five LMs on Alpaca vs. full fine-tuning.
@@
Reducing overfitting in meta-training: Recall that in the meta-training stage, we apply multitask training to the combined samples of all the tasks. We conduct a preliminary experiment where we use a sharpness-reduced training algorithm (i.e., sharpness-aware minimization Foret et al. (2021); Zhang et al. (2024)) for multitask training (in place of SGD). We find that this can reduce the approximation error (relative to full fine-tuning results) by 23%. Further applying multitask and meta-learning techniques to improve GradEx would be an interesting question for future work.
@@ 4.3 Downstream evaluation for fine-tuning
Baselines: We compare our algorithms to forward selection (FS) and random ensemble (RE) with full fine-tuning. Additionally, we compare with four baseline methods, including fine-tuning a single model on all tasks with LoRA (MTL), Data Selection with Importance Resampling (DSIR) Xie et al. (2023), Data-Efficient Fine-Tuning with cross-task nearest neighbors clustering based on feature similarity (DEFT) Ivison et al. (2023), and Low-rank gradiEnt Similarity Search (LESS) Xia et al. (2024). For each baseline, the performance is reported based on fine-tuning a pretrained LM with the same number of trainable parameters on the selected subset of tasks.
Evaluation metrics: For instruction tuning, we follow the protocol of Wang et al. (2023a) to evaluate the toxicity and truthfulness of fine-tuned models in ToxiGen Hartvigsen et al. (2022) and TruthfulQA Lin et al. (2022), respectively. For ToxiGen, we measure the percentage of toxic outputs generated by the model. For TruthfulQA, we evaluate model accuracy for identifying truthful statements. For StrategyQA and CommonsenseQA, we measure the accuracy of the generated answers.
@@ 4.3.1 Results for instruction tuning
First, we compare GradEx-FS with directly testing the pretrained LM and fine-tuning the last output layer of the LM (Feature Transfer). We find that GradEx-FS outperforms both methods by 9% and 11% on average, respectively. Compared to fine-tuning a model on all tasks (MTL), GradEx-FS leads to an 8% improvement.
@@ 4.3.2 Results for chain-of-thought fine-tuning
First, GradEx-FS outperforms directly testing the pretrained model and feature transfer by 7% and 16%, respectively. GradEx-FS also improves the accuracy by 5% over MTL. Second, even with the estimation, we see similar downstream performance to forward selection and random ensemble (with true fine-tuning) while using only 3% and 0.5% of the total computation costs. Lastly, both GradEx-FS and GradEx-RE can outperform the four baseline selection methods by 2.4% and 3.5% (on average), respectively.
As for data selection, we first cluster the samples into $n$ groups. We vary $n$ between $50$, $100$, $200$, and $400$. We also evaluate our approach without the clustering step. We find that using $100$ groups yields the best performance and indeed outperforms not using clustering by 1.4%. Therefore, we set the number of groups as $100$ in the chain-of-though fine-tuning.
@@ 5 Related Work
Data modeling: There is a growing line of work on understanding the role of individual samples in deep networks and large models. Kozareva et al. (2011) propose a centrality and regularization-based algorithm by leveraging graph structures to explore the relationship between multiple instances in a semantic network. Wettig et al. (2024) select pre-training data by training a rater model to evaluate four data quality criteria that align with human intuitions. Thrush et al. (2024) select data based on the losses and perplexities of existing LLMs on pretraining texts. In contrast, this work focuses on task-relatedness in fine-tuning language models. This is a simpler problem than pre-training, as the number of tasks is smaller. A potential adaptation could be splitting pre-training into multiple phases and applying our techniques to develop a curriculum for pre-training.
Our approach is also related to the work of Ilyas et al. (2022) and Park et al. (2023) on using random sampling for data modeling and attribution. The difference is that we utilize the linearization property of LLMs for fine-tuning close to the initialization. A closely related work by Li et al. (2023b) introduces a surrogate modeling approach to identify subsets of source tasks that benefit a target task. They show that a linear surrogate model is particularly helpful for identifying negative transfers, including higher-order negative transfers from a subset of source tasks to the target task. This higher-order transfer corresponds to a boosting procedure that turns out to be very helpful in practice Li et al. (2023a). In addition, this surrogate modeling approach can also be used in data augmentation for finding tree-structured compositions Li et al. (2024a). It would be very interesting to understand the underlying mathematical structure of LLMs that enables such geometric properties in practice Zhang (2023).
@@ 6 Conclusion
Acknowlegement: Thanks to the anonymous referees and the action editor for their constructive feedback. The work of D. Li, Z. Zhang, and H. Zhang is partly supported by NSF award IIS-2412008. D. Li was also partially funded by a PhD fellowship from JPMorgan Chase & Co. Any views or opinions expressed herein are solely those of the authors listed, and may differ from the views and opinions expressed by JPMorgan Chase & Co. or its Affiliates.
@@ Appendix C Ommited Materials for Section 3
0.05% $6_{\pm 0.6}\times 10^{-4}$ $6_{\pm 0.5}\times 10^{-5}$ $8_{\pm 1.1}\times 10^{-4}$ $3_{\pm 0.4}\times 10^{-5}$
0.10% $1_{\pm 0.2}\times 10^{-3}$ $3_{\pm 0.3}\times 10^{-4}$ $2_{\pm 0.4}\times 10^{-3}$ $1_{\pm 0.3}\times 10^{-4}$
0.15% $2_{\pm 0.1}\times 10^{-3}$ $6_{\pm 0.6}\times 10^{-4}$ $3_{\pm 0.4}\times 10^{-4}$ $3_{\pm 0.7}\times 10^{-4}$
0.20% $3_{\pm 0.3}\times 10^{-3}$ $1_{\pm 0.2}\times 10^{-3}$ $3_{\pm 0.7}\times 10^{-4}$ $4_{\pm 0.9}\times 10^{-4}$
0.25% $4_{\pm 0.2}\times 10^{-2}$ $5_{\pm 0.8}\times 10^{-3}$ $7_{\pm 0.6}\times 10^{-2}$ $5_{\pm 0.8}\times 10^{-3}$
@@ D.3 Comparison of GPU hours
# FLOPs ($\downarrow$) ToxiGen TruthfulQA CommonsenseQA StrategyQA
@@ D.5 Complete pseudocode for subset selection procedures
Input: The training datasets of $n$ data sources
Require: A multitask fine-tuning algorithm $f$
@@
2:for $i=1,2,\ldots,n$ do
3:     for each task not selected in current subset $j\in[n]/S_{\text{cur}}$  do
4:         Add task $j$ to current subset: $S_{j}\leftarrow S_{\text{cur}}\bigcup\{j\}$
5:         Estimate $\hat{f}(S_{j})$ using GradEx
@@
7:     Choose the task with the highest value function $j^{\star}\leftarrow\arg\max_{j\in[n]/S_{\text{cur}}}\hat{f}(S_{j})$
8:     Add the dataset to the current subset: $S_{\text{cur}}\leftarrow S_{\text{cur}}+j^{\star}$, if $\hat{f}(S_{j})<\hat{f}(S_{\text{cur}})$; otherwise, break
@@
Parameter: The size of each sampled subset $\alpha$, and the number of sampled subsets $m$; a selection threshold $\gamma$
@@
3:     Estimate $\hat{f}(S_{k})$ using GradEx
@@
5:Estimate a score for each dataset $i$ as the value function averaged over subsets that include $i$:
@@
6:Select a subset of tasks by thresholding the scores: $S^{\star}=\{i\ |\ T_{i}<\gamma,\ \forall\ i=1,2,\ldots,n\}$